OpenStudy (anonymous):

multiply and simplify: x+3 * x+4 ----- ------------ x x^2+7x+12

OpenStudy (ivettef365):

do you know how to solve x^2+7x+12 ??

OpenStudy (anonymous):

noo :/

OpenStudy (ivettef365):

ok, you need to find a multiple of 12 that when you add = 7 ?

OpenStudy (ivettef365):

that would be 4 x 3 = 12 and 4 + 3 = 7

OpenStudy (anonymous):

ohh okay isee

OpenStudy (ivettef365):

so to solve that would be (x+4)(x+3)

OpenStudy (anonymous):

okayy now what

OpenStudy (anonymous):

x^2+7x+12 =X^2+3x+4x+12 =x(x+3)+4(x+3) =(x+4)(x+3)

OpenStudy (ivettef365):

so now you have x+3 x+4 ----- . -------- x (x+3)(x+4)

OpenStudy (anonymous):

okay

OpenStudy (ivettef365):

now you can cancel and what do you have left ?

OpenStudy (anonymous):

what do i cancel

OpenStudy (anonymous):

{(x+3)/x}*{x+4/(x+4)(x+3)} x is ur answer after all cancellation

OpenStudy (anonymous):

so how would i write the final answer

OpenStudy (ivettef365):

what is the same from top and bottom that is like if I have 4/4 that would be = 1 actually you are not cancelling you are converting to 1

OpenStudy (anonymous):

ohhhh

OpenStudy (anonymous):

so how would i write the final answer

OpenStudy (anonymous):

cause what that other guys said confused me lol

OpenStudy (ivettef365):

you have left the following 1 1 ------- . ------- = ? just multiply and what do you get ? x 1

OpenStudy (anonymous):

1 ---- 1x

OpenStudy (ivettef365):

correct, which is the same as 1/x

OpenStudy (anonymous):

okay, thank you for teaching me i kinda understand it now :)

OpenStudy (ivettef365):

yw :)